### v1

# Exploring the In-Context Learning Capabilities of LLMs for Money Laundering Detection in Financial Graphs

Abstract
The complexity and interconnectivity of entities involved in money laundering demand investigative reasoning over graph-structured data. This paper explores the use of large language models (LLMs) as reasoning engines over localized subgraphs extracted from a financial knowledge graph. We propose a lightweight pipeline that retrieves k-hop neighborhoods around entities of interest, serializes them into structured text, and prompts an LLM via few-shot in-context learning to assess suspiciousness and generate justifications. Using synthetic anti-money laundering (AML) scenarios that reflect common laundering behaviors, we show that LLMs can emulate analyst-style logic, highlight red flags, and provide coherent explanations. While this study is exploratory, it illustrates the potential of LLM-based graph reasoning in AML and lays groundwork for explainable, language-driven financial crime analytics.
123
1 Introduction
As financial data becomes increasingly interconnected, institutions are turning to knowledge graphs to capture the complex relationships between customers, accounts, transactions, devices, and geographic entities. These graphs provide a rich structure for reasoning about risk, behavior, and anomalies. However, leveraging their full potential requires tools capable of understanding and explaining multi-hop, heterogeneous patterns.
Recent advances in large language models (LLMs) have demonstrated their ability to reason over structured data, including relational chains and graph-like abstractions. A growing body of research has shown that LLMs can be adapted to perform graph reasoning tasks via graph-structured prompting [4], or language-guided message passing [9]. Surveys on this topic highlight a promising direction for unifying natural language and graph reasoning under a shared framework [8].
Building on this momentum, we investigate whether LLMs can support investigative-style reasoning over financial knowledge graphs. These graphs represent real-world financial systems, where relational complexity is not only common but essential to risk analysis, trend analysis, etc., [5]. Our goal is to evaluate the feasibility of using LLMs as interpretable reasoning engines over such data — particularly in domains where explainability, transparency, and analyst collaboration are critical.
As a motivating case study, we focus on anti-money laundering (AML) in this study. AML is a domain where identifying suspicious behavior involves tracing intricate financial flows across individuals, shell entities, and intermediary accounts [2, 3]. AML and financial crime investigations increasingly rely on graph-based representations to detect patterns such as layering, rapid movement of funds, and shared infrastructure [7, 6].
For this study, we utilize the publicly available IBM AML Synthetic Dataset [1], which simulates financial activities through a realistic knowledge graph of accounts, banks, and transactions, designed to model common laundering typologies based on real-world intelligence. This allows us to explore LLM-based reasoning in a controlled yet representative environment.
This work positions LLMs not merely as text generators, but as in-context reasoning engines capable of interpreting graph-structured financial activity. In-context learning offers a flexible, parameter-free mechanism to inject domain knowledge and reasoning capabilities into black-box LLMs, aligning naturally with investigative workflows. By leveraging this mechanism, we evaluate whether LLMs can emulate analyst-style reasoning in AML detection without retraining or labeled supervision.
We propose a lightweight reasoning pipeline in which:
1.
A -hop subgraph around a transaction or entity of interest is extracted from a financial knowledge graph stored in a graph database,
2.
The subgraph is serialized into a structured textual format designed to preserve topology and semantics,
3.
An LLM is prompted in a few-shot setup to assess suspiciousness and provide an investigator-style justification.
We evaluate this pipeline using synthetic AML scenarios that reflect realistic laundering behaviors, constructed from publicly available datasets. Our early results indicate that LLMs can highlight meaningful red flags, trace multi-step patterns, and generate rationales that align with human intuition. This case study illustrates the potential of LLMs for reasoning over financial knowledge graphs, offering a path toward explainable, language-driven fraud analytics. To our knowledge, this is the first work to apply in-context learning with LLMs for money laundering detection over financial graphs. The proposed pipeline bridges localized graph structure with natural language reasoning in a scalable and interpretable way.
2 Methodology
We propose a three-stage framework for enabling LLM-based reasoning over financial knowledge graphs: (1) extracting subgraphs around suspicious transaction edges, (2) converting them into a structured textual format, and (3) prompting an LLM in a few-shot setup informed by money laundering typologies. The model is tasked not only with predicting whether the transaction is suspicious but also providing a rationale and identifying any typologies that apply, which allows us to assess its reasoning transparency and hallucination rate.
2.1 Financial Graph Representation and Subgraph Extraction
Money laundering schemes often involve complex transactional relationships that can be naturally represented as graphs. In this work, we use the publicly available IBM AML Synthetic Dataset [1], which models financial activities as a knowledge graph.
In this graph, nodes represent accounts (individual or corporate) and banks, while edges represent financial transactions between accounts. Each node is annotated with metadata such as entity type and account creation date; each transaction edge is labeled with attributes including transaction amount and timestamp.
Our goal is to reason about the suspiciousness of specific transactions by analyzing the surrounding graph structure. For each candidate transaction (i.e., a transfer between two accounts), we extract a localized -hop subgraph centered on the transaction edge. The subgraph includes the source and destination accounts, their neighboring transactions within hops, and any associated bank nodes.
An example of an extracted subgraph is shown below. Each extracted structure provides the LLM with a focused local context to assess whether the transaction may reflect suspicious activity.
This extraction strategy frames the detection task as an edge-centric classification problem over the transaction graph, where the objective is to predict whether the given transaction edge represents money laundering activity.
2.2 Textual Serialization
The extracted subgraph is then serialized into a structured textual representation suitable for LLM input. Each node is expressed as a role-tagged entity, and each edge is rendered as a natural language relation, preserving directionality, timestamps, and values. The textual format preserves the topological structure of the graph, laying out interactions in logical order to facilitate multi-hop reasoning. No further summarization or feature engineering is applied.
To illustrate this process, we include a representative serialization from an actual extracted subgraph depicted in Figure 2 and used in our prompting pipeline:
2.3 Few-Shot Prompting with AML Typologies
A defining characteristic of anti-money laundering (AML) investigations is the presence of recurring structural patterns in transactional behavior. These patterns have been extensively documented in financial intelligence reports and academic studies, including the IBM AML dataset, which models them as prototypical graph topologies. Figure 3 visualizes eight canonical laundering structures used in our prompting framework: (a) fan-out, (b) fan-in, (c) gather-scatter, (d) scatter-gather, (e) simple cycle, (f) random, (g) bipartite, and (h) stack.
To help the LLM learn to reason over financial subgraphs, our approach uses a few-shot prompting setup where each demonstration example showcases one of these laundering patterns. Each example consists of (i) a brief serialized subgraph in the textual format described in Section 2.2, and (ii) a natural language rationale that explains why the activity is suspicious, grounded in the structure of the graph. Both the graph serialization and the rationale are derived from samples in the IBM AML dataset, which provides well-annotated synthetic cases.
The prompt includes eight such examples, each corresponding to one of the patterns shown in Figure 3, before appending the test case. By explicitly conditioning the LLM on real-world-style typologies and human-readable rationales, we align the reasoning process with investigative workflows. Importantly, these demonstrations emphasize graph topology and multi-hop interactions rather than superficial statistical cues, which makes the model more likely to generalize beyond seen entities.
This methodology leverages the core advantage of few-shot learning via in-context learning: enabling the model to acquire structural priors without any fine-tuning. It effectively bootstraps the LLM into a role akin to a fraud analyst by showing what suspicious behavior “looks like” in multiple formats. The choice to use real pattern types, realistic transactions, and natural language justifications makes the prompting setup robust, interpretable, and well-aligned with human investigator expectations.
To further refine the model’s discrimination ability, we also include several examples of non-suspicious financial activity—subgraphs that contain no red flags—within the same prompt. This ensures that the model learns not only to detect suspicious patterns, but also to distinguish them from legitimate transactional behavior.
The final structure of the LLM prompt follows the template below, showing how demonstration subgraphs and rationales are interleaved. The subgraphs are serialized using the format shown previously in the Serialized Subgraph Example box in Section 2.2.
3 Experiments and Results
We evaluate the proposed reasoning pipeline on the IBM AML Synthetic Dataset, which simulates real-world money laundering typologies through graph-structured financial transactions. Our primary goal at this stage is to assess both the classification ability and the reasoning quality of a large language model (LLM) over structured financial subgraphs.
3.1 Experimental Setup
We construct a balanced evaluation set consisting of 2,000 examples: 1,000 laundering transactions and 1,000 non-laundering transactions. While this does not reflect the class imbalance typically observed in real-world AML investigations, it enables a controlled evaluation of the model’s ability to distinguish suspicious patterns without being dominated by prior class frequencies. This design isolates the core reasoning capabilities of the LLM pipeline and ensures that performance metrics reflect meaningful classification behavior rather than dataset skew. Extending the evaluation to realistic, highly imbalanced distributions remains an important direction for future work.
Each subgraph is serialized following the format described in Section 2.2 and embedded into the final prompt template described in Section 2.3. The prompt includes eight suspicious pattern examples and four non-suspicious examples, with task instructions asking the model to: (1) classify the subgraph as suspicious or not suspicious, and (2) list any observed money laundering patterns. We use the GPT-4o model accessed via the OpenAI API (temperature = 0). Each test sample is independently prompted with no carryover memory.
3.2 Sample Model Output
Below we present a representative output generated by the LLM for a test subgraph classified as suspicious:
3.3 Classification Performance
Given the computational cost of LLM inference at scale, rerunning the model over multiple randomized test sets was infeasible. Instead, to assess the robustness of our results without incurring additional inference cost, we apply bootstrap resampling over the fixed set of 2,000 model outputs. Specifically, we generate 1,000 bootstrap samples (sampling with replacement) and recompute key evaluation metrics for each resample. This procedure allows us to estimate the mean and 95% confidence intervals for classification accuracy, precision, recall, and the F1 scores. Reporting confidence intervals provides insight into the stability of both classification and reasoning performance. Table 1 summarizes the model’s binary classification performance.
| Metric | Value (95% CI) |
| Overall Accuracy | 63.7% 3.3% |
| Precision | 62.7% 4.4% |
| Recall | 67.7% 4.4% |
| F1 Score | 65.1% 3.9% |
3.4 Reasoning and Pattern Recognition Performance
To evaluate the model’s reasoning quality and tendency to hallucinate patterns, we measure the precision and recall of the patterns listed by the LLM relative to ground-truth pattern labels provided by the IBM dataset.
| Pattern | Precision (95% CI) | Recall (95% CI) |
| Fan-out | 75.1% 4.0% | 80.3% 3.8% |
| Fan-in | 73.8% 4.1% | 81.5% 4.0% |
| Gather-scatter | 66.2% 4.5% | 70.7% 4.3% |
| Scatter-gather | 64.9% 4.5% | 68.8% 4.3% |
| Simple cycle | 62.5% 4.6% | 65.9% 4.5% |
| Random | 51.5% 5.0% | 53.5% 4.9% |
| Bipartite | 55.0% 5.0% | 56.1% 5.0% |
| Stack | 52.0% 5.1% | 58.0% 5.0% |
Detection performance generally reflects the underlying structural complexity of the laundering patterns. Highly organized patterns such as fan-out and fan-in achieve the strongest results, while moderately complex patterns like gather-scatter, scatter-gather, and simple cycle show intermediate performance. Among the harder patterns, such as random, bipartite, and stack, we observe some variability, reflecting the increased difficulty in consistently capturing less structured behaviors. Overall, these preliminary findings suggest that prompting LLMs over serialized financial graphs provides a viable pathway for both suspicious activity classification and explainable pattern-driven reasoning.
4 Conclusion and Future Work
In this paper, we explored the feasibility of using large language models (LLMs) to perform structured reasoning over financial knowledge graphs for anti-money laundering (AML) investigations. By constructing localized subgraphs around transactions, serializing them into textual prompts, and conditioning the LLM with a few-shot setup based on canonical laundering patterns, we demonstrated that LLMs can not only classify suspicious behavior but also provide interpretable rationales for their decisions through in-context learning.
While this work presents an initial exploration rather than a production-ready system, it is not intended as a fully benchmarked replacement for existing graph-based fraud detection methods. Instead, our goal was to assess the feasibility and interpretability of in-context LLM reasoning over financial knowledge graphs. This study provides a conceptual foundation for hybrid approaches that integrate explainability and investigative logic, while also recognizing that computational and scalability challenges must be addressed for future practical deployments.
One of the most important advantages of this approach is its strong interpretability. Unlike traditional graph neural networks (GNNs) or other embedding-based methods, which often act as black boxes, the LLM produces human-readable explanations that align with investigative reasoning. This opens the door for models that are not only accurate but also transparent and auditable, addressing a key pain point in the deployment of machine learning systems for financial crime detection.
While promising, the approach also presents practical limitations. In particular, the computational cost of prompting large language models at inference time makes it impractical to apply as a stand-alone solution for large-scale production environments. Instead, we envision a hybrid system in which a lightweight classifier, such as a GNN or a graph feature engineering model combined with a gradient boosting classifier, performs initial triage. Only cases with low classification confidence (borderline decisions) would be passed to the LLM for additional reasoning. This selective invocation strategy would allow the system to maintain scalability while leveraging LLM-generated rationales to enhance decision quality and explainability. In future work, we plan to benchmark such hybrid systems and evaluate the impact of LLM-based reasoning on overall detection performance and model transparency compared to existing graph-based methods.
In addition to hybridizing with traditional graph learning methods, we also plan to extend this line of research by investigating more advanced GNN+LLM architectures that better integrate structural reasoning into the model itself. Furthermore, we intend to study whether smaller LLMs, fine-tuned specifically on financial graph reasoning tasks, can achieve similar or better interpretability with significantly reduced computational overhead. These directions will help unlock scalable, interpretable, and highly effective money laundering and fraud detection systems for real-world financial environments.
Disclaimer
This material is provided for informational purposes only and is not intended to be investment advice or a recommendation to take any particular investment action.